OpenStudy (anonymous):

Leah's house has a media room with a large flat-screen television to watch movies. The media room television has a width of 8 feet and a height of 4.5 feet. The flat-screen television in Leah's living room has a height of 27 inches and has the same width to height ratio as the media room television. What is the width of Leah's living room television? A. 48 inches B. 15 inches C. 32 inches D. 64 inches

OpenStudy (anonymous):

@Isaiah.Feynman

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@niksva

OpenStudy (anonymous):

@kingofme

OpenStudy (anonymous):

@Drsuz98

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@WeepingAngel123

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

@Abdulhameed

OpenStudy (anonymous):

c is dat right @kingofme

OpenStudy (anonymous):

@RoseDryer

OpenStudy (anonymous):

IS IT C

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yep c i kno

OpenStudy (anonymous):

i thought it was b

OpenStudy (destinymasha):

Okay, ratio = they are related in some way 4.5 /8 = 27/ ? 4.5 divided by 27 = 6, so multiply 6 x 8 =48 SO 4.5/8 = 27/48 -->check that by dividing each numerator by 6 I think 48 should be the answer